# Non-equilibrium Quantum Monte Carlo Algorithm for Stabilizer Rényi Entropy in Spin Systems

### v3

@@ II Methods
$M_{n}(\rho)=\frac{1}{1-n}\ln\sum_{\sigma\in\mathcal{P}_{N}}\frac{1}{2^{N}}\operatorname{Tr}[\rho\sigma]^{2n}-S_{n}(\rho)$ (1)
@@
Figure 1: Tensor-network diagram for partition function $\mathcal{Z}_{B}$ (see Eqn. (4), $n=2$ in this figure). The tensor products of replicas $e^{-\beta\hat{H}}/2$ are represented as stacks of $e^{-\beta\hat{H}}/2$ tensors. $\bm{a},\bm{b},...,\bm{i},\bm{j},...,\bm{a}^{\prime},\bm{b}^{\prime},...,\bm{i}^{\prime},\bm{j}^{\prime},...\in\{0,1\}^{\otimes N}$ are $N$-spin product states labeling the legs of tensors. $T(B)$ is represented as a $4nN$ tensor, where the green (red) circles denote $B_{s}=0$ ($B_{s}=1$) for site $s=1,2,\cdots,N$. For $T>0$, the PBC in imaginary time direction impose the constraints $\bm{a}=\bm{a}^{\prime},\bm{b}=\bm{b}^{\prime},\cdots$, while for $T=0$ case $\bm{a},\bm{b},...,\bm{a}^{\prime},\bm{b}^{\prime},...$ have OBC but are required to be contracted with $|\phi\rangle$, e.g., $\langle\bm{a}|\phi\rangle$. Regarding the QMC sampling for $\mathcal{Z}(B)$, $\hat{Q}=\{\hat{Q}^{[r]}_{1,2},r=1,2,\cdots,2n\}$ denote the operator lists sampled from $e^{-\beta\hat{H}/2}$ in each replica.
@@
For example, the first term in Eqn. (1) for $T>0$ states can be written as
$M_{n}(\rho)=\frac{1}{1-n}\ln\frac{\mathcal{Z}_{[N]}}{\mathcal{Z}_{\varnothing}}$ (3)
where we define $\mathcal{Z}_{B}$ for any subset of spins $B$ from the whole system as
@@
and
$T(B)=\frac{1}{2^{N_{B}}}\otimes_{s=1}^{N}\left(\mathcal{T}^{B_{s}}I_{2n}^{1-B_{s}}\right)$ (5)
$B_{s}=1$ if the spin at site $s\in B$, and 00 otherwise, $N_{B}$ is the total number of spins in $B$ and we define in Eqn (6) the connection tensor $\mathcal{T}$, a $4n$-leg tensor connecting spins of the same site indices from the $2n$ replicas of $\operatorname{Tr}[e^{-\beta H}]$ and $I_{2n}$ is the identity tensor with the same shape as $\mathcal{T}$.
$\mathcal{T}^{j_{1},j_{2},\cdots,j_{2n}}_{k_{1},k_{2},\cdots,k_{2n}}=\otimes_{r=1}^{2n}I_{k_{r}}^{j_{r}}+\otimes_{r=1}^{2n}Z_{k_{r}}^{j_{r}}+\otimes_{r=1}^{2n}X_{k_{r}}^{j_{r}}+\otimes_{r=1}^{2n}Y_{k_{r}}^{j_{r}}$ (6)
To apply the non-equilibrium QMC algorithm, which amounts to sampling the evolution from $\mathcal{Z}_{\varnothing}$ and $\mathcal{Z}_{[N]}$, we introduce a parametrized partition function interpolating these two through a weight function $g(\lambda,N_{B})\equiv\lambda^{N_{B}}(1-\lambda)^{N-N_{B}}$ [40]:
$\mathcal{Z}(\lambda)=\sum_{B\subseteq[N]}\lambda^{N_{B}}(1-\lambda)^{N-N_{B}}\mathcal{Z}_{B},\quad\lambda\in[0,1]$ (7)
Henceforth, we can rewrite SRE (Eqn. (3)) in an integral form:
@@
Our final step is thus to use QMC to sample these paths and evaluate the work $W$ from each path. This is done through random walkers where we slowly increase $\lambda$ from 0 to 1, and record the work $W$ as the increment of $\Delta\ln g$. Each random walker samples configurations from an extended configuration space $\varkappa\equiv\{\bm{a},\bm{b},\cdots\bm{a^{\prime}},\bm{b^{\prime}},\hat{Q},B\}$ which labels the legs of the tensors $\exp(-\beta\hat{H}/2)$ (lowercase letters $\bm{a},\bm{b},\cdots\bm{a^{\prime}},\bm{b^{\prime}}$), the internal operator lists ($\hat{Q})$ needed to sample from each replica of $\exp(-\beta\hat{H}/2)$ and the set of connected sites ($B$) (see Fig. 1). The configurations are sampled with probability
@@
$\begin{split}\text{Wt}(\varkappa=\{\bm{a},\bm{b},\cdots,\hat{Q},B\};\lambda)=&g(\lambda,N_{B})\langle\bm{a}|\hat{Q}^{[1]}_{1}|\bm{i}\rangle\langle\bm{b}|\hat{Q}^{[2]}_{1}|\bm{j}\rangle\cdots\\ &T(\bm{i},\bm{j},\cdots;\bm{i}^{\prime},\bm{j}^{\prime},\cdots|B)\\ &\langle\bm{i}^{\prime}|\hat{Q}^{[1]}_{2}|\bm{a}^{\prime}\rangle\langle\bm{j}^{\prime}|\hat{Q}^{[2]}_{2}|\bm{b}^{\prime}\rangle\cdots\end{split}$ (10)
where
$T(\bm{i},\bm{j},\cdots;\bm{i}^{\prime},\bm{j}^{\prime},\cdots|B)=\frac{1}{2^{N_{B}}}\prod_{s\in B}\mathcal{T}^{i_{s},j_{s},\cdots}_{i^{\prime}_{s},j^{\prime}_{s},\cdots}=1$ (11)
@@
for spin states $\bm{i},\bm{j},\cdots;\bm{i}^{\prime},\bm{j}^{\prime},\cdots$ corresponding to non-zero elements of $\mathcal{T}$, and 0 otherwise. To get the samples from $\lambda=0$ which are the initial configurations of each random walker, we can use standard QMC on each replica $\exp(-\beta\hat{H}/2)$.
This new framework we’ve developed which represents the ratio of partition functions as tensor networks and then evaluates that ratio using non-equilibrium QMC can be generalized to various other quantities including other types of Rényi entropy calculations, e.g., entanglement Rényi negativity (ERN), a measure of quantum correlation within mixed states [45, 46, 47, 48, 49, 50], and Shannon Rényi entropy, also known as participation Rényi entropy (PRE), a measure for the wave function localization in the computational basis [51, 52, 53, 54]. We refer to the supplemental materials (SM) for discussion of these generalizations.
III Numerical Experiments
$\hat{H}=-J\sum_{\langle i,j\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}$ (12)
where SSE is applied to update the configurations [55, 56, 57, 58].
In Fig. 2, we show the comparisons between the 2nd SRE density $m_{2}(\rho)\equiv M_{2}(\rho)/N$ from QMC and those from matrix product states (MPS) in various scenarios: (a) ground state SRE of 1D Ising ring (i.e., periodic boundary conditions, PBC); (b) ground state SRE of 2D Ising cylinder (i.e., PBC on one spatial direction and open boundary condition on the other, PBC-OBC); (c) finite-$T$ SRE of 2D Ising cylinder. They all display quantitative agreements when comparing against with tensor-network methods. Remarkably, in case (c), the noticeable discrepancies between the QMC and MPS data result from the errors of limited bond dimension $\chi$ and the finite Trotter splitting time interval $\tau$ for the MPS method as can be seen from the observation that the MPS data converges towards the QMC data as we increase $\chi$ and decrease $\tau$. This suggests that to reach the same accuracy as QMC, the MPS-based methods requires more computational resources and careful extrapolation. In SM, we also compare our results with those from Ref. [27].
In Fig. 3, we report our results on larger systems, and we are able to draw the following conclusions: (1) 2nd SRE mainly grows linearly with system size for short-range correlated pure states, as indicated by the ground states on 2D transverse field Ising model. (2) At $T>0$, there exist subleading terms from 2nd SRE weakly dependent on $T$, as evident from the crossing in Fig. 3(b). In Fig. 4, we show the 2nd SRE versus a broader range of temperature, and find two extremals at $\beta\simeq 0.36$ and $0.78$ for sufficiently large systems.
IV Computational Cost
For $T=0$ or ground state SRE, we use projector QMC and the key difference for the time complexity is that the length of the operator list $m$ has to be sufficiently large to decrease the contributions from the excited states and guarantee the convergence to the ground state. We argue that the lower bound for $m$ such that the relative bias error to the exact ground state SRE $M_{n}(g)$ is below $\delta_{r}$ is given by (see SM for the derivation)
V Conclusion
In this work, we develop a novel algorithm to compute the SRE of finite temperature and ground state systems in all dimensions for sign-problem free Hamiltonians. To accomplish this, we show how to contract a tensor network representation of the SRE using non-equilibrium QMC. It has proven to be accurate and efficient, regardless of the spatial dimensions and temperature of the model, as observed from the numerical experiments on the transverse-field Ising model and the complexity analysis of this algorithm. We expect this algorithm would serve as a useful numerical tool to understand quantum systems from the new perspective of quantum magic. Interesting future directions include, but are not limited to, an in-depth study of phase transition features from $n(\geq 2)$th SRE consisting of the derivatives [45], the application to models with deconfined quantum critical points [60, 61, 62, 63, 64] or field theory [21, 65, 66], the extension of this approach to Hamiltonians with sign problems, and the exploration of replacing the non-equilibrium QMC algorithm with other efficient algorithms for evaluating the exponential observable [37, 38, 35],.
Acknowledgements.
We would like to thank T.C. Lu, Z.Y. Meng, J. Zhao, J. D’Emidio, M. Dalmonte, E. Tirrito, M. Frau and P. Tarabunga for insightful discussions and valuable comments. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and which is supported by funds from the University of Illinois at Urbana-Champaign.
Appendix A Non-equilibrium QMC Algorithm
In this section, we provides some details and remarks on the non-equilibrium QMC algorithm.
Note that we have transformed the summation over exponentially many Pauli strings into a concise partition function. This transformation is transparent from the following identities:
$\begin{split}\mathcal{Z}_{[N]}&=\sum_{\sigma\in\mathcal{P}_{N}}\frac{1}{2^{N}}\operatorname{Tr}[e^{-\beta\hat{H}}\sigma]^{2n}\\ &=\operatorname{Tr}\left[\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}\left(\frac{1}{2^{N}}\sum_{\sigma\in\mathcal{P}_{N}}\sigma^{\otimes 2n}\right)\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}\right]\\ &=\operatorname{Tr}\left[\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}T([N])\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}\right]\end{split}$ (14)
Essentially, the summation over the exponential number of terms is greatly simplified by summing over each site individually first.
For the connection tensor, when $n=2$, only 16 out of the 256 elements of $\mathcal{T}$ are non-zero (and equal to 2), such that the sampling space is well constrained and the sign problem will not occur from the presence of $\mathcal{T}$.
However, one can always view this tensor network contraction as a sum over terms which can then be sampled by Monte Carlo. A naive approach to this sampling is to sample terms from $\mathcal{Z}_{\varnothing}$ and then count how many of these terms appear in $\mathcal{Z}_{N}$ [29, 33]. This can be done in the standard QMC fashion where one samples operator lists for $e^{-\beta\hat{H}/2}$ [55, 56, 57]. Unfortunately, for the natural case where the SRE grows linearly with system size, the ratio $\mathcal{Z}_{[N]}/\mathcal{Z}_{\varnothing}$ will get exponentially small, and one will only exponentially rarely see a term which simultaneously belongs to $\mathcal{Z}_{[N]}$ and $\mathcal{Z}_{\varnothing}$ resulting in an exponential computational cost [29, 33]. This is the exponential observable problem encountered in the ERE computation and it has been solved recently by sampling reweighted observable with polynomial complexity [40, 41, 42, 34, 37, 38, 35, 36].
The basic implementation of the non-equilibrium QMC algorithm for finite-$T$ systems is as follows (also see Fig. 5):
@@ Appendix A Non-equilibrium QMC Algorithm
Moreover, we note two parallelization strategies to speed up this algorithm: (1) Divide the evolution of $\lambda$ into $K$ smaller intervals: $[0,1]=[0,1/K]\cup[1/K,2/K]\cup\cdots\cup[1-1/K,1]$; Apply the above algorithm to each interval independently; Then the original SRE is given by the sum from these smaller intervals [41]. (2) As each side walker from Step (3) evolves independently, one can distribute each of them to different processors and only collect the observable $W$ from them at the end of the evolution.
Figure 5: Illustration of the non-equilibrium QMC algorithm. (a) SRE is viewed as the ratio between two partition functions in the partition function space, and the QMC algorithm essentially sample the paths between these two partition functions. (b) Diagram for the implementation steps of the algorithm.
Appendix B Implementation of SSE on Ising Model - Finite Temperature
In this section, we describe how to apply our algorithm for computing the SRE using the SSE algorithm as the QMC on the transverse field Ising Model.  [55, 56, 57, 58].
@@ Appendix B Implementation of SSE on Ising Model - Finite Temperature
There are two stages of configuration updates within SSE: diagonal update and off-diagonal update. The diagonal update here is the identical to the standard one [55]. For the off-diagonal update, we use the cluster updating scheme, where we group the legs from all the operators into clusters, each of them containing the whole two-spin operators but ending at single-spin operators. After all clusters are formed, we choose to flip each cluster with $1/2$ probability following the Swendsen-Wang scheme [58].
When a cluster reaches one of the $2n$ legs from $\mathcal{T}$, we use the multi-branch clustering to continue growing the cluster out of $\mathcal{T}$ [57]. More specifically, the cluster will branch out from the $n-1$ legs of $\mathcal{T}$ on the same side as the reached leg, so that the cluster enters into other replicas. This ensures that after the update, the spin configuration still corresponds to a non-zero entry of $\mathcal{T}$.
Appendix C Implementation of SSE on Ising Model - Zero Temperature
To obtain the zero-temperature or the ground state SRE on the Ising model, we use projector QMC with the product state $|00\cdots 0\rangle$ as the initial trial state $|\phi\rangle$ and an operator list with long enough length $m$. Different from the finite-$T$ algorithm, here there are open boundary conditions on the imaginary axis, that is, the cluster growth also ends at the boundary. In the zero temperature situation, we fix the initial product state of $|00...0\rangle$ at $\bm{a},\bm{b},\bm{c},\bm{d}$ and $\bm{a^{\prime}},\bm{b^{\prime}},\bm{c^{\prime}},\bm{d^{\prime}}$. Note that this gives the SRE of the ground state which has overlap with the $|00...0\rangle$ sector.

### v4

@@ II Methods
$\begin{split}\widetilde{M}_{n}(\rho)=&\frac{1}{1-n}\ln\sum_{\sigma\in\mathcal{P}_{N}}\frac{1}{2^{N}}\operatorname{Tr}[\rho\sigma]^{2n}-S_{n}(\rho)\\ \equiv&M_{n}(\rho)-S_{n}(\rho)\end{split}$ (1)
@@
Figure 1: Tensor-network diagram for the partition function $\mathcal{Z}_{B}$ (see Eqn. (9)) with $42n=$ 4 replicas. Each (roman indexed) line is a spin in $\{0,1\}$ and gray boxes represent the tensor $e^{-\beta\hat{H}/2}$. $T(B)$ is represented as a $4nN$ tensor consisting of a tensor product of $\mathcal{T}$’s where $B_{s}=1$, each of which are represented as connected red dots spanning replicas over their respective site $s$. All other sites have $B_{s}=0$ corresponding to identities (green dots). Note that of the drawn sites, $B_{s}=1$ on sites 2 and $N$ and $B_{s}=0$ on site 1. For finite temperature, the PBC in the imaginary time direction imposes the constraints $\bm{a}=\bm{a}^{\prime},\bm{b}=\bm{b}^{\prime},\cdots$, while for the zero temperature case, the sites $\bm{a},\bm{b},...,\bm{a}^{\prime},\bm{b}^{\prime},...$ are contracted with $|\phi\rangle$, e.g., $\langle\bm{a}|\phi\rangle$. When this tensor network is sampled with QMC, all spins are sampled and the operator list $\hat{Q}=\{\hat{Q}^{[r]}_{1,2},r=1,2,\cdots,2n\}$ is sampled from each gray box.
@@
The first term in Eqn. (1) for $T>0$ states can be written (combining Eqn. (1) and (2)) in terms of the ratio of two partition functions:
$\begin{split}M_{n}(\rho)=&\frac{1}{1-n}\ln\frac{\sum_{\sigma\in\mathcal{P}_{N}}\frac{1}{2^{N}}\operatorname{Tr}[e^{-\beta\hat{H}}\sigma]^{2n}}{\left(\operatorname{Tr}e^{-\beta\hat{H}}\right)^{2n}}\\ \equiv&\frac{1}{1-n}\ln\frac{\mathcal{Z}_{[N]}}{\mathcal{Z}_{\varnothing}}\end{split}$ (3)
Writing
$\begin{split}\mathcal{Z}_{[N]}&=\operatorname{Tr}\left[\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}\left(\frac{1}{2^{N}}\sum_{\sigma\in\mathcal{P}_{N}}\sigma^{\otimes 2n}\right)\left(e^{-\beta\hat{H}/2}\right)^{\otimes 2n}\right]\end{split}$ (4)
we can then rewrite the summation over exponentially many Pauli strings as
$\begin{split}\sum_{\sigma\in\mathcal{P}_{N}}\sigma^{\otimes 2n}&=\otimes_{s=1}^{N}\left(\sum_{\sigma_{s}\in\{I,Z,X,Y\}}\sigma_{s}^{\otimes 2n}\right)\\ &\equiv\otimes_{s=1}^{N}\mathcal{T}\end{split}$ (5)
where the connection tensor $\mathcal{T}$ is a $4n$-leg tensor connecting spins of the same site from each of the $2n$ replicas, whose tensor elements are
$\mathcal{T}^{j_{1},j_{2},\cdots,j_{2n}}_{k_{1},k_{2},\cdots,k_{2n}}=\prod_{r=1}^{2n}I_{j_{r},k_{r}}+\prod_{r=1}^{2n}Z_{j_{r},k_{r}}+\prod_{r=1}^{2n}X_{j_{r},k_{r}}+\prod_{r=1}^{2n}Y_{j_{r},k_{r}}$ (6)
and which can be represented with a tensor-network diagram as
[Uncaptioned image] (7)
Using the notation that $\mathcal{T}^{0}=I^{\otimes 2n}$ is the identity for each of the $2n$ replicas, we can define the tensor
$T(B)=\frac{1}{2^{N_{B}}}\otimes_{s=1}^{N}\mathcal{T}^{B_{s}}$ (8)
where $B$ is a set of sites (to be thought of as the sites where are connected over the replicas by $\mathcal{T}$) and $B_{s}=1$ if site $s\in B$ and 00 otherwise, $N_{B}=|B|$ is the total number of sites in $B$. This lets us write
@@
and define both $\mathcal{Z}_{[N]}$ and $\mathcal{Z}_{\varnothing}$ where $[N]$ is the set of all sites and $\varnothing$ is the empty set $\mathcal{Z}_{B}$ for sets of intermediate size will be useful in developing an algorithm to compute the ratio $\mathcal{Z}_{\varnothing}/\mathcal{Z}_{[N]}$.
However, one can always view this tensor network contraction as a sum over terms which can then be sampled by Monte Carlo. A naive approach to this sampling is to sample terms from $\mathcal{Z}_{N}$ and then count how many of these terms appear in $\mathcal{Z}_{\varnothing}$ [29, 33]. This can be done in the standard QMC fashion where one samples operator lists for $e^{-\beta\hat{H}/2}$ [44, 45, 46]. Unfortunately, for the natural case where the SRE grows linearly with system size, the ratio $\mathcal{Z}_{\varnothing}/\mathcal{Z}_{[N]}$ will get exponentially small, and one will only exponentially rarely see a term which simultaneously belongs to $\mathcal{Z}_{[N]}$ and $\mathcal{Z}_{\varnothing}$ resulting in an exponential computational cost [29, 33]. This is the exponential observable problem encountered in the ERE computation and it has been solved recently by sampling reweighted observable with polynomial complexity [40, 41, 42, 34, 37, 38, 35, 36].
We apply the non-equilibrium QMC algorithm to evaluate the ratio $\mathcal{Z}_{\varnothing}/\mathcal{Z}_{[N]}$ efficiently, which amounts to sampling a set of paths which evolve from $\mathcal{Z}_{\varnothing}$ and $\mathcal{Z}_{[N]}$. To this end, we introduce a parametrized partition function interpolating these two partition functions through a weight function $g(\lambda,N_{B})\equiv\lambda^{N_{B}}(1-\lambda)^{N-N_{B}}$ [40]:
$\mathcal{Z}(\lambda)=\sum_{B\subseteq[N]}g(\lambda,N_{B})\mathcal{Z}_{B},\quad\lambda\in[0,1]$ (10)
Henceforth, we can rewrite Eq. (3) in an integral form:
@@
Our final step is thus to use QMC to sample these paths and evaluate the work $W$ from each path. This is done through random walkers where we slowly increase $\lambda$ from 0 to 1, and record the work $W$ as the increment of $\Delta\ln g$. Each random walker samples configurations from an extended configuration space $\varkappa\equiv\{\bm{a},\bm{b},\cdots\bm{a^{\prime}},\bm{b^{\prime}},\hat{Q},B\}$ which include the legs of the tensors $\exp(-\beta\hat{H}/2)$ (lowercase letters $\bm{a},\bm{b},\cdots\bm{a^{\prime}},\bm{b^{\prime}}$), the internal operator lists ($\hat{Q})$ needed to sample from each replica of $\exp(-\beta\hat{H}/2)$ and the set of connected sites ($B$) (see Fig. 1). The configurations are sampled with probability
@@
$\begin{split}\text{Wt}(\varkappa=\{\bm{a},\bm{b},\cdots,\hat{Q},B\};\lambda)=&g(\lambda,N_{B})\langle\bm{a}|\hat{Q}^{[1]}_{1}|\bm{i}\rangle\langle\bm{b}|\hat{Q}^{[2]}_{1}|\bm{j}\rangle\cdots\\ &T(B)^{\bm{i},\bm{j},\cdots}_{\bm{i}^{\prime},\bm{j}^{\prime},\cdots}\\ &\langle\bm{i}^{\prime}|\hat{Q}^{[1]}_{2}|\bm{a}^{\prime}\rangle\langle\bm{j}^{\prime}|\hat{Q}^{[2]}_{2}|\bm{b}^{\prime}\rangle\cdots\end{split}$ (13)
where the tensor elements of $T(B)$ are
$T(B)^{\bm{i},\bm{j},\cdots}_{\bm{i}^{\prime},\bm{j}^{\prime},\cdots}=\frac{1}{2^{N_{B}}}\prod_{s\in B}\mathcal{T}^{i_{s},j_{s},\cdots}_{i^{\prime}_{s},j^{\prime}_{s},\cdots};$ (14)
@@
note that these tensor elements are either 1 if the spin states $\bm{i},\bm{j},\cdots;\bm{i}^{\prime},\bm{j}^{\prime},\cdots$ correspond to non-zero elements for each of $\mathcal{T}$, or 0 otherwise. When $n=2$, for example, only 16 out of the 256 elements of $\mathcal{T}$ are non-zero (and equal to 2), such that the sampling space is well constrained and the sign problem will not be induced by the use of $\mathcal{T}$. To get the samples from $\lambda=0$ which are the initial configurations of each random walker, we can use standard QMC on each replica $\exp(-\beta\hat{H}/2)$.
The basic implementation of the non-equilibrium QMC algorithm for finite-$T$ systems is as follows:
@@
Moreover, we note two parallelization strategies to speed up this algorithm: (1) Divide the evolution of $\lambda$ into $K$ smaller intervals: $[0,1]=[0,1/K]\cup[1/K,2/K]\cup\cdots\cup[1-1/K,1]$; Apply the above algorithm to each interval independently (e.g., $t_{i}=k/K$, $t_{f}=(k+1)/K$); Then the original SRE is given by the sum from these smaller intervals [41]. (2) As each side walker from Step (3) evolves independently, one can distribute each of them to different processors and only collect the observable $W$ from them at the end of the evolution.
This new framework we have developed which represents the ratio of partition functions as tensor networks and then evaluates that ratio using non-equilibrium QMC can be generalized to various other quantities including other types of Rényi entropy calculations, e.g., entanglement Rényi negativity (ERN), a measure of quantum correlation within mixed states [48, 49, 50, 51, 52, 53], and Shannon Rényi entropy, also known as participation Rényi entropy (PRE), a measure for the wave function localization in the computational basis [54, 55, 56, 57]. We refer to the Appendix. E for discussion of these generalizations.
III Numerical Experiments
$\hat{H}=-J\sum_{\langle i,j\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}$ (16)
where stochastic series expansion (SSE) is applied to update the configurations [44, 45, 46, 58].
III.1 Implementation of SSE on Ising Model - Finite Temperature
In this section, we describe how SSE works for the non-equilibrium algorithm on the transverse field Ising Model  [44, 45, 46, 58].
@@ III.1 Implementation of SSE on Ising Model - Finite Temperature
There are two stages of configuration updates within SSE: diagonal update and off-diagonal update. The diagonal update here is identical to the standard one [44]. For the off-diagonal update, we use the cluster updating scheme, where we group the legs from all the operators into clusters, each of them containing the whole two-spin operators but ending at single-spin operators. After all clusters are formed, we choose to flip each cluster with $1/2$ probability following the Swendsen-Wang scheme [58].
When a cluster reaches one of the $2n$ legs from $\mathcal{T}$, we use the multi-branch clustering to continue growing the cluster out of $\mathcal{T}$ [46]. More specifically, the cluster will branch out from the $n-1$ legs of $\mathcal{T}$ on the same side as the reached leg, so that the cluster enters into other replicas. This ensures that after the update, the spin configuration still corresponds to a non-zero entry of $\mathcal{T}$. For example, as shown in Fig. 2(a), when a cluster from the second replica reaches one leg of $\mathcal{T}$ (indicated by the blue arrow going towards $\mathcal{T}$), the cluster grows into the rest of the replicas (as indicated by the green arrows going outwards from $\mathcal{T}$). Similarly, in Fig. 2(b), we also show the diagram for the directed loop update scheme if we apply the algorithm to models such as XXZ model, where we let the directed loop from the second replica grow into the first replica, and a new directed loop form within third and forth replicas.
Figure 2: Schematic diagrams for (a) cluster updating and (b) loop updating when the cluster or the directed loop reaches the leg of $\mathcal{T}$.
III.2 Implementation of SSE on Ising Model - Zero Temperature
To obtain the zero-temperature or the ground state SRE on the Ising model, we use projector QMC with the product state $|00\cdots 0\rangle$ as the initial trial state $|\phi\rangle$ and an operator list with long enough length $m$. Different from the finite-$T$ algorithm, here there are open boundary conditions on the imaginary time axis, that is, the cluster growth also ends at the boundary. In the zero temperature situation, we fix the initial product state of $|00...0\rangle$ at $\bm{a},\bm{b},\bm{c},\bm{d}$ and $\bm{a^{\prime}},\bm{b^{\prime}},\bm{c^{\prime}},\bm{d^{\prime}}$. Note that this gives the SRE of the ground state which has overlap with the $|00...0\rangle$ sector.